1.    What concentration of NaF do you need to add to a 0.15 M HF solution to get a buffer solution pH of 9.5? The Ka for HF is 7.4x10-4.


1
Expert's answer
2022-05-10T06:21:09-0400

pKa = - log (7.4×10-4) = 3.13

pH = pKa + log ([NaF]/[HF]) 

9.50 = 3.13 + log ([NaF]/[HF]) 

log ([NaF]/[HF]) = 6.37

[NaF] / 0.15 = 106.37

[NaF] = 106.37/0.15

Most likely the problem has an incorrect pH value, usually it is not alkaline.